# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v3

@@ 3.4 When $\delta\ll s$
$\theta=2\pi n-\frac{8\pi\delta}{\beta s\tanh\left(\frac{\pi^{2}}{\beta}\right)}+O(t^{-1}),$ (3.36)
where $n=0$ or $1$, depending on its monodromic behavior (n=1n=-1 is not possible as the permissible range of $\theta$ is $2\pi<\theta<2\pi$). It turns out, as can be seen in figure 6, 1z1-z revolves around the origin once in the anticlockwise direction as it evolves with time. Therefore 1z1-z in the large-$t$ limit does not simply approach 1, but $e^{2\pi i}$ instead; the correct choice of $n$ would be $n=1$.
@@
As usual in the AdS/CFT, this holographic CFT calculation is equivalently described by the gravitational dual of AdS3. As sketched in figure 9, the local operator and the mixed-state excitation are dual to a deficit angle and a localized black hole in AdS3. The analysis of holographic local operator quenches Nozaki:2013wia ; Bhattacharyya:2019ifi shows that their cores are located at $z=\sqrt{\delta^{2}+(T+t)^{2}}$ and $z=\sqrt{s^{2}+(T+t-t_{M})^{2}}$ under time evolution, working with the Poincaré AdS3 metric $ds^{2}=z^{-2}(dz^{2}-dT^{2}+dx^{2})$. The propagation of the entangled pair created by the local operator is dual to that of the gravitational shockwave produced by the time-dependent deficit angle. We expect that some of this shockwave is absorbed by the localized black hole, dual to the mixed-state quench. This provides a gravitational interpretation of entanglement suppression.
@@
$\theta=2\pi n+\frac{8\pi^{2}\delta}{\beta^{2}\sinh^{2}\left(\frac{\pi^{2}}{\beta}\right)t}+O(t^{-2}),$ (3.45)
where n=1n=-1 or 0. Similarly for $\bar{z}$, we get
@@
$\bar{\theta}=2\pi\bar{n}-\frac{\delta}{t_{M}}+O(t^{-1}),$ (3.47)
where $\bar{n}=0$ or $1$. The only permissible pairs of $(n,\bar{n})$ are $(0,1)$ and (1,0)(-1,0), so $\Delta S_{A^{c}}$ comes out to be
@@ 3.5 When $s\ll\delta$
The fact that 1z1-z only makes a $\pi$ revolution has significant implications when we consider a case II setup. When $a<x_{P}$, 1z1-z loses time dependence and is always $\theta=\pm\pi$ where its sign depends on the channel. This means that it is impossible to set $\theta=0$ or even $|\theta|\ll 1$ at $t=0$, which is necessary in order to impose $\Delta S_{A}=0$ at $t=0$. So, necessarily $\Delta S_{A}\neq 0$ at $t=0$, which is highly counterintuitive from the point of view of causality. This may be due to the fact that the excited state is a mixed state, which is a point that we also touched upon in the previous subsection when considering $\delta\ll s$. Another possibility is that since $s\ll\delta$, the local operator may not be as localized, potentially giving rise to nonzero $\Delta S_{A}$ even at early times.
@@ 3.7 Energy density
Figure 15: Energy density of a local operator quench with (blue) and without (orange) a mixed-state excitation in the background. The excitations are located at $x_{P}=-1,t_{M}=1$. Other fixed parameters are taken to be $s=1,\beta=1$. For the left panel, $\delta=0.15$, and for the right panel, $\delta=0.20$.
@@ 4.1 Validity of our approximation
Figure 16: Plots of $\Delta S_{A}$ as $\delta/s$ is varied. On the left we have $\delta/s<1$, and on the right we have $\delta/s>1$. The subsystem is taken to be $a=0,b=1000$ and the separation between excitations is taken to be lightlike and small relative to the subsystem ($x_{P}=-1,t_{M}=1$). $\beta=1$ (BTZ phase). This is a snapshot of $\Delta S_{A}$ when $t=500$. Details of computation for $\delta/s>1$ regarding the choice of branch is discussed in section 4.3.
This singular behavior of $\Delta S_{A}$ clearly looks unphysical. Moreover, we saw earlier in section 3.6 that, even away from $\delta/s=O(1)$, the analytic expressions obtained in the limits $\delta/s\ll 1$ and $\delta/s\gg 1$ for the total entanglement entropy $S_{A}$, when the separation is lightlike, do not map to each other under the duality transformation (3.63). If we remember how we computed the entanglement entropy, we notice that the only approximation we employed which requires proper justification is to regard correlation functions on a torus as those on an infinite cylinder by picking up the BTZ saddle. Even though this is a quite common approximation in AdS/CFT, we have to be careful now because we inserted a heavy local operator in addition to the mixed-state excitation. The latter is described by the BTZ black hole after a coordinate transformation. The former excitation introduces a time-evolving deficit angle, which produces a huge backreaction on the BTZ geometry.
Though our HHLL approximation is still able to incorporate perturbative backreactions Fitzpatrick:2014vua , we may expect non-perturbative effects such as other saddle contributions, which we did not take into account. This non-perturbative contribution becomes enhanced when the two excitations approach each other. We note that the two excitations are separated by $|s-\delta|$ in the Euclidean time direction even when they collide in the real spacetime. Therefore, it is natural to think that we can ignore non-perturbative interactions between the two excitations when $\delta/s\ll 1$ or $\delta/s\gg 1$. This heuristic holographic argument is depicted in figure 9. For a more rigorous justification, we would need to calculate the full correlation function on a torus and compare it with that on an infinite cylinder. Since the former is currently not well-understood, we shall leave this as a future problem.
@@ 4.2 When $\delta<s$
for integer $n$ at lightlike separation. The full dependence of $\theta$ on $\delta/s$ is depicted in figure 17.
We can see that the once $\delta$ crosses over $\tanh\left(\frac{n\beta}{4}\right)s$ for odd $n$, the monodromic behavior $(n,\bar{n})\mapsto(n+1,\bar{n})$ (as was the case for $\delta\ll s$) shifts to $(n,\bar{n})\mapsto(n-1,\bar{n})$. We should also note that away from (4.5), $|\theta-2\pi n|\ll 1$ can no longer be satisfied, so, as commented in section 3.5, $\Delta S_{A}=0$ at $t=0$ is unattainable for case II setups.
@@ 4.3 When $s<\delta$
Figure 18: Plots of $\zeta_{\mathcal{O}^{\dagger}}^{\beta/2\pi i}=e^{w_{\mathcal{O}^{\dagger}}}$ as $\delta$ is varied, when $\delta\ll s$ (left) and $s\ll\delta$ (right).
@@
Figure 19: Plots of $\Delta S_{A}$ as $t_{M}$ is varied, when the positive branch is chosen (left) and the negative branch is chosen (right). The subsystem is taken to be $a=10,b=20$ and the operator is inserted at $x_{P}=-1$. The snapshot of $\Delta S_{A}$ is taken at $t=16$. The ratios of the regulators are chosen to be $s/\delta=0.01,0.50,0.99$ going from top to bottom, where $\delta$ is fixed to be $1$. $\beta=1$ (BTZ phase) for all plots.
@@ 5 Conclusion and discussion
Our analysis of entanglement entropy of doubly-excited states provides a new approach to the scattering problem of strongly-coupled quantum field theories. Our entanglement suppression is due to strong interactions in two-dimensional holographic CFTs. On the other hand, it is known that in integrable CFTs, entanglement suppression does not occur. In this way, entanglement suppression is expected to be a new probe of the chaotic dynamics in quantum field theories or even in more general quantum many-body systems. It would be intriguing to investigate entanglement suppression in various strongly-correlated quantum many-body systems. At the same time, it would also be intriguing to understand the basic mechanism of why entanglement suppression occurs from the viewpoint of quantum many-body dynamics. One useful approach is to use the connection between holographic spacetimes and tensor networks Swingle:2009bg ; Nozaki:2012zj ; Qi:2013caa ; Pastawski:2015qua ; Hayden:2016cfa ; Milsted:2018san . In particular the idea of path integral optimization Caputa:2017urj ; Caputa:2017yrh will provide a direct connection between our Euclidean path integral construction and its tensor network description. We hope to come back to this problem in the near future.

### v4

@@ 3.4 When $\delta\ll s$
$\theta=2\pi m-\frac{8\pi\delta}{\beta s\tanh\left(\frac{\pi^{2}}{\beta}\right)}+O(t^{-1}),$ (3.36)
where $m=0$ or $1$, depending on its monodromic behavior (m=1m=-1 is not possible as the permissible range of $\theta$ is $2\pi<\theta<2\pi$). It turns out, as can be seen in figure 6, 1z1-z revolves around the origin once in the anticlockwise direction as it evolves with time. Therefore 1z1-z in the large-$t$ limit does not simply approach 1, but $e^{2\pi i}$ instead; the correct choice of $n$ would be $m=1$.
@@
As usual in the AdS/CFT, this holographic CFT calculation is equivalently described by the gravitational dual of AdS3. As sketched in figure 9, the local operator and the mixed-state excitation are dual to a deficit angle and a localized black hole in AdS3. The analysis of holographic local operator quenches Nozaki:2013wia; Bhattacharyya:2019ifi shows that their cores are located at $z=\sqrt{\delta^{2}+t^{2}}$ and $z=\sqrt{s^{2}+(t-t_{M})^{2}}$ respectively under time evolution, working with the Poincaré AdS3 metric $ds^{2}=z^{-2}(dz^{2}-dt^{2}+dx^{2})$. The propagation of the entangled pair created by the local operator is dual to that of the gravitational shockwave produced by the time-dependent deficit angle. We expect that some of this shockwave is absorbed by the localized black hole, dual to the mixed-state quench. This provides a gravitational interpretation of entanglement suppression.
@@
$\theta=2\pi m+\frac{8\pi^{2}\delta}{\beta^{2}\sinh^{2}\left(\frac{\pi^{2}}{\beta}\right)t}+O(t^{-2}),$ (3.45)
where m=1m=-1 or 0. Similarly for $\bar{z}$, we get
@@
$\bar{\theta}=2\pi\bar{m}-\frac{\delta}{t_{M}}+O(t^{-1}),$ (3.47)
where $\bar{m}=0$ or $1$. The only permissible pairs of $(m,\bar{m})$ are $(0,1)$ and (1,0)(-1,0), so $\Delta S_{A^{c}}$ comes out to be
@@ 3.5 When $s\ll\delta$
The fact that 1z1-z only makes a $\pi$ revolution has significant implications when we consider a case II setup. When $a<x_{P}$, 1z1-z loses time dependence and is always $\theta=\pm\pi$ where its sign depends on the channel. This means that it is impossible to set $\theta=0$ or even $|\theta|\ll 1$ for $t<x_{P}-a$, which is necessary in order to impose $\Delta S_{A}=0$ at early times. So, necessarily $\Delta S_{A}\neq 0$ for $t<x_{P}-a$, which is highly counterintuitive from the point of view of causality.
This conundrum can be partially salvaged by remembering that these local quenches are inserted by hand, so the non-vanishing of $\Delta S_{A}$ at early times is actually contained within $t_{M}<t<x_{P}-a$. At $t<t_{M}$, the mixed-state local quench has not yet taken place, so we should revert to results from section 2 and conclude that $\Delta S_{A}=0$ at early times before $t=t_{M}$.
The non-vanishing of $\Delta S_{A}$ for $t_{M}<t<x_{P}-a$ remains a mystery. It may be due to the fact that the background on which the local operator quench takes place is a mixed state, and that information about this background somehow spills over to $\Delta S_{A}$. This is a point that was touched upon in the previous subsection when considering $\delta\ll s$, as we saw that $\Delta S_{A}\neq\Delta S_{A^{c}}$. Another possibility is that the cylindrical approximation of the setup geometry — and therefore the HHLL approximation of conformal blocks — fails at lightlike separation of excitations. As we will see in the following subsection where we test how duality (3.20) holds up in the setup of a double local quench, we obtain an anomalous term at lightlike separation, which suggests a possible breakdown of the approximation. A further discussion on the validity of the HHLL approximation is provided in section 4.1.
@@ 3.7 Energy density
Figure 15: Energy density of a local operator quench with (blue) and without (orange) a mixed-state excitation in the background. The excitations are located at $x_{P}=-1,t_{M}=1$. The snapshots are taken at $t=5$ and other fixed parameters are taken to be $s=1,\beta=1$. For the left panel, $\delta=0.15$, and for the right panel, $\delta=0.20$.
@@ 4.1 Validity of our approximation
Figure 16: Plots of $\Delta S_{A}$ as $\delta/s$ is varied. On the left we have $\delta/s<1$, and on the right we have $\delta/s>1$. The subsystem is taken to be $a=0,b=1000$ and the separation between excitations is taken to be lightlike and small relative to the subsystem ($x_{P}=-1,t_{M}=1$). $\beta=1$ (BTZ phase). These are snapshots of $\Delta S_{A}$ when $t=500$. Details of computation for $\delta/s>1$ regarding the choice of branch is discussed in section 4.3.
@@
This singular behavior of $\Delta S_{A}$ clearly looks unphysical. Moreover, we saw earlier in section 3.6 that, even away from $\delta/s=O(1)$, the analytic expressions obtained in the limits $\delta/s\ll 1$ and $\delta/s\gg 1$ for the total entanglement entropy $S_{A}$, when the separation is lightlike, do not map to each other under the duality transformation (3.63). If we remember how we computed the entanglement entropy, we notice that the only approximation we employed which requires proper justification is to regard correlation functions on a torus as those on an infinite cylinder by picking up the BTZ saddle. Even though this is a quite common approximation in AdS/CFT, we have to be careful now because we inserted a heavy local operator in addition to the mixed-state excitation. The latter is described as a BTZ black hole after a coordinate transformation. The former excitation introduces a time-evolving deficit angle, which produces a huge backreaction on the BTZ geometry.
Though our HHLL approximation is still able to incorporate perturbative backreactions Fitzpatrick:2014vua, we may expect nonperturbative effects such as other saddle contributions, which we did not take into account. This nonperturbative contribution becomes enhanced when the two excitations approach each other. Noting that the two excitations are separated in the original Euclidean geometry by $(X_{\text{sep}},\bar{X}_{\text{sep}})=(i(s-\delta)+x_{P}+t_{M},-i(s-\delta)+x_{P}-t_{M})$, we can reasonably expect nonperturbative interactions between the two excitations to be insignificant when $\delta/s\ll 1$, $\delta/s\gg 1$, and/or $|x_{P}+t_{M}|\gg\delta,s$. This heuristic holographic argument is depicted in figure 9. For a more rigorous justification, we would need to calculate the full correlation function on a torus and compare it to that on an infinite cylinder. Since the former is currently not well-understood, we shall leave this as a future problem.
@@ 4.2 When $\delta<s$
for integer $n$ at lightlike separation. The full dependence of $\theta$ on $\delta/s$ is depicted in figure 17. We can see that the once $\delta$ crosses over $\tanh\left(\frac{n\beta}{4}\right)s$ for odd $n$, the monodromic behavior $(m,\bar{m})\mapsto(m+1,\bar{m})$ (as was the case for $\delta\ll s$) shifts to $(m,\bar{m})\mapsto(m-1,\bar{m})$. We should also note that away from (4.5), $|\theta-2\pi m|\ll 1$ can no longer be satisfied, so, as commented in section 3.5, $\Delta S_{A}=0$ at $t=0$ is unattainable for case II setups.
@@ 4.3 When $s<\delta$
Figure 18: Contours of $\zeta_{\mathcal{O}^{\dagger}}^{\beta/2\pi i}=e^{w_{\mathcal{O}^{\dagger}}}$ as $\delta$ is varied, when $\delta\ll s$ (left) and $s\ll\delta$ (right).
@@
Figure 19: Plots of $\Delta S_{A}$ as $t_{M}$ is varied, when the positive branch is chosen (left) and the negative branch is chosen (right). The subsystem is taken to be $a=10,b=20$ and the operator is inserted at $x_{P}=-1$. These snapshots of $\Delta S_{A}$ are taken at $t=16$. The ratios of the regulators are chosen to be $s/\delta=0.01,0.50,0.99$ going from top to bottom, where $\delta$ is fixed to be $1$. $\beta=1$ (BTZ phase) for all plots.
@@ 5 Conclusion and discussion
Our analysis of entanglement entropy of doubly-excited states provides a new approach to the scattering problem of strongly-coupled quantum field theories. Our entanglement suppression is due to strong interactions in two-dimensional holographic CFTs. On the other hand, it is known that in integrable CFTs, entanglement suppression does not occur. In this way, entanglement suppression is expected to be a new probe of the chaotic dynamics in quantum field theories or even in more general quantum many-body systems. It would be intriguing to investigate entanglement suppression in various strongly-correlated quantum many-body systems. At the same time, it would also be intriguing to understand the basic mechanism of why entanglement suppression occurs from the viewpoint of quantum many-body dynamics. One useful approach is to use the connection between holographic spacetimes and tensor networks Swingle:2009bg; Nozaki:2012zj; Qi:2013caa; Pastawski:2015qua; Hayden:2016cfa; Milsted:2018san. In particular, the idea of path integral optimization Caputa:2017urj; Caputa:2017yrh will provide a direct connection between our Euclidean path integral construction and its tensor network description. We hope to come back to this problem in the near future.